# Dynamic and Distributed Routing in IoT Networks based on Multi-Objective Q-Learning

### v1

@@ VI-C1 Overall Reward
We present both episodic rewards (figure 6(a)) and accumulated returns (figure 6(b)) of the proposed Distributed DPQ-Learning routing algorithm against the six baseline routing algorithms. From figure 6(a), we can observe that the proposed Distributed DPQ-Learning routing algorithm learns to optimize the overall return by doing more exploration than exploitation in the first $1000$ episodes. The pareto front is also learned in this phase. In contrast to the sequential exploration-exploitation scheme, the simultaneous scheme involves both exploration and exploration occurring hand-in-hand while gradually reducing exploration and increasing exploitation. From figure 6(a), we notice the performance of the proposed Distributed DPQ-Learning routing algorithm may not be superior in the beginning phases of exploration, but after a few exploration ($1000$ episodes), it outperforms all the baselines with respect to episodic rewards. The proposed protocol also dynamically and immediately adjusts to changing preferences after a few explorations, while other baselines fail to do so. For example, SMORLR learns to adapt after each preference change. However, the process is slow, and the preferences change before it can adapt to match the proposed protocol’s performance. Further, R2LTO, RLBEEP, FMOLD, and RLBR consistently focus on a fixed objective preference, mostly prioritizing energy efficiency. They tend to learn certain intervals to increase returns, but they still fall short of the proposed method by 1213×12-13$\times$12 - 13 × approximately. RSPAR performs, on average, 344% less efficiently than the proposed approach. Thus, while certain baseline methods may perform well under specific preferences, the proposed algorithm consistently outperforms them and adapts quickly to changing preferences.
@@
Figure 6(b) illustrates cumulative performance over time. Initially, all (baseline and proposed) algorithms start with zero cumulative returns. At the beginning, RSPAR and RLBR outperform the proposed DPQ-Learning algorithm. However, after a few explorations, the proposed approach starts utilizing the learned Pareto front and consistently outperforms the baselines, with the performance gap widening as more episodes are completed. Figure 6(b) shows that after $10,000$ episodes, the accumulated return of the proposed method surpasses SMORLR by $111.11\%$, RSPAR by $322.22\%$, R2LTO and RLBEEP by $375\%$, FMOLD by $850\%$, and RLBR by $927\%$ approximately. Similarly, significant improvements were seen in Experiment $4$, where preferences were randomly altered in each episode. Figure 6(c) shows that the proposed routing protocol consistently achieves significantly higher accumulated returns than the baseline, with the performance gap growing as episodes progress. By the end of $5000$ episodes, the proposed algorithm achieves an accumulated return of around $1750$, which is 5.48×5.48$\times$5.48 × that of RSPAR and RLBEEP, 8.21×8.21$\times$8.21 × that of SMORLR, 34×34$\times$34 × that of FMOLD, and 36×36$\times$36 × that of R2LTO and RLBR. Based on these results, we observe that the proposed method consistently outperforms all baselines in terms of overall returns and adapts rapidly to changes in preferences regardless of exploration policy.
VI-C2 Energy
@@
In this subsection, we assess the energy-efficiency of the proposed and baseline routing algorithms on individual objectives when preferences between them change periodically and frequently. Figure 7(a) highlights the episodic energy consumption (in mJ) across various routing algorithms over $10,000$ episodes, where the preference changes after every $1000$ episode. On average, the proposed distributed DPQ-Learning routing protocol demonstrates $65-80$% lower energy consumption than the other routing protocols. The proposed approach consumed the minimum energy whenever a higher preference was given to the energy objective (intervals $80001000-2000,4000-5000,and6000-$ 8000) while performing reasonably well even when the preference for the energy objective was lower. On average, SMORLR consumes approximately $80\%$ more energy than the proposed method, although it gradually converges closer to it with each preference shift. However, the more frequently preferences change, the less efficient SMORLR becomes. R2LTO, RLBEEP, and RLBR consume around $60\%$ more energy than the proposed method. Since they focus primarily on energy considerations, they perform the highest among baselines. However, the proposed approach outperforms them because of a more efficient reward formulation. FMOLD consumes around $160\%$ more energy than the proposed method while maintaining a consistent performance. RSPAR exhibits the energy inefficiency $\approx 200\%$ more energy consumption than the proposed algorithm.
@@ VI-C3 Packet Delivery
Figure 8: Packet Delivery rate in the Simultaneous exploration-exploitation scheme.
@@
VII Conclusions
This paper presents a dynamic and distributed IoT routing protocol based on multi-objective Q-learning, designed to adapt to changing priorities in real-time. This approach optimizes key metrics like overall reward, energy efficiency and PDR. By leveraging distributed learning at individual nodes, the method avoids the need for centralized control, enhancing scalability and robustness. Simulations show that this approach outperforms existing algorithms in adaptability, energy efficiency, and packet delivery. The proposed approach shows consistent out-performance irrespective of the exploration strategies, and preference change patterns. The study highlights the potential of reinforcement learning for dynamic IoT networks and suggests future work on extending the algorithm to address security and more complex objectives. This makes it a promising solution for real-time, flexible routing in evolving IoT environments.

### v2

@@ VII-C1 Overall Reward
We present both episodic rewards (figure 6(a)) and accumulated returns (figure 6(b)) of the proposed Distributed DPQ-Learning routing algorithm against the six baseline routing algorithms. From figure 6(a), we can observe that the proposed Distributed DPQ-Learning routing algorithm learns to optimize the overall return by doing more exploration than exploitation in the first $1000$ episodes. The pareto front is also learned in this phase. In contrast to the sequential exploration-exploitation scheme, the simultaneous scheme involves both exploration and exploration occurring hand-in-hand while gradually reducing exploration and increasing exploitation. From figure 6(a), we notice the performance of the proposed Distributed DPQ-Learning routing algorithm may not be superior in the beginning phases of exploration, but after a few exploration ($1000$ episodes), it outperforms all the baselines with respect to episodic rewards. The proposed protocol also dynamically and immediately adjusts to changing preferences after a few explorations, while other baselines fail to do so. For example, SMORLR learns to adapt after each preference change. However, the process is slow, and the preferences change before it can adapt to match the proposed protocol’s performance. Further, R2LTO, RLBEEP, FMOLD, and RLBR consistently focus on a fixed objective preference, mostly prioritizing energy efficiency. They tend to learn certain intervals to increase returns, but they still fall short of the proposed method by $12-13\times$ approximately. RSPAR performs, on average, 344% less efficiently than the proposed approach. Thus, while certain baseline methods may perform well under specific preferences, the proposed algorithm consistently outperforms them and adapts quickly to changing preferences.
@@
Figure 6(b) illustrates cumulative performance over time. Initially, all (baseline and proposed) algorithms start with zero cumulative returns. At the beginning, RSPAR and RLBR outperform the proposed DPQ-Learning algorithm. However, after a few explorations, the proposed approach starts utilizing the learned Pareto front and consistently outperforms the baselines, with the performance gap widening as more episodes are completed. Figure 6(b) shows that after $10,000$ episodes, the accumulated return of the proposed method surpasses SMORLR by $111.11\%$, RSPAR by $322.22\%$, R2LTO and RLBEEP by $375\%$, FMOLD by $850\%$, and RLBR by $927\%$ approximately. Similarly, significant improvements were seen in Experiment $4$, where preferences were randomly altered in each episode. Figure 6(c) shows that the proposed routing protocol consistently achieves significantly higher accumulated returns than the baseline, with the performance gap growing as episodes progress. By the end of $5000$ episodes, the proposed algorithm achieves an accumulated return of around $1750$, which is $5.48\times$ that of RSPAR and RLBEEP, $8.21\times$ that of SMORLR, $34\times$ that of FMOLD, and $36\times$ that of R2LTO and RLBR. Based on these results, we observe that the proposed method consistently outperforms all baselines in terms of overall returns and adapts rapidly to changes in preferences regardless of exploration policy.
VII-C2 Energy
@@
In this subsection, we assess the energy-efficiency of the proposed and baseline routing algorithms on individual objectives when preferences between them change periodically and frequently. Figure 7(a) highlights the episodic energy consumption (in mJ) across various routing algorithms over $10,000$ episodes, where the preference changes after every $1000$ episode. On average, the proposed distributed DPQ-Learning routing protocol demonstrates 658065-80% lower energy consumption than the other routing protocols. The proposed approach consumed the minimum energy whenever a higher preference was given to the energy objective (intervals 10002000,400050001000-2000,4000-5000, and 600080006000-8000) while performing reasonably well even when the preference for the energy objective was lower. On average, SMORLR consumes approximately $80\%$ more energy than the proposed method, although it gradually converges closer to it with each preference shift. However, the more frequently preferences change, the less efficient SMORLR becomes. R2LTO, RLBEEP, and RLBR consume around $60\%$ more energy than the proposed method. Since they focus primarily on energy considerations, they perform the highest among baselines. However, the proposed approach outperforms them because of a more efficient reward formulation. FMOLD consumes around $160\%$ more energy than the proposed method while maintaining a consistent performance. RSPAR exhibits the energy inefficiency $\approx 200\%$ more energy consumption than the proposed algorithm.
@@ VII-C3 Packet Delivery
Figure 8: Packet Delivery Ratio in the Simultaneous exploration-exploitation scheme.
@@
VIII Conclusions
This paper presents a distributed multi-objective Q-learning–based routing algorithm for IoT and wireless sensor networks. The method learns multiple per-preference Q-tables in parallel and uses a greedy interpolation policy to adapt instantly to changing preference weights without retraining or central coordination. Theoretical analysis establishes a uniform near-optimality guarantee, explaining the stability and effectiveness of the proposed interpolation mechanism. Simulation results confirm consistent gains over existing schemes, demonstrating that the proposed Dynamic Preference Q-Learning framework provides a light-weight and theoretically grounded approach for real-time distributed routing in dynamic IoT environments.
Shubham Vaishnav is currently pursuing Ph.D. at the Department of Computer and Systems Science, Stockholm University, Sweden. This research program is funded by Digital Futures. He received Bachelors’ and Masters’ degrees in Computer Science and Engineering from the Indian Institute of Technology (Indian School of Mines), Dhanbad. His research interests include Multiobjective and Adaptive decision-making in IoT and Edge networks, Reinforcement learning, Federated Learning, and data-driven decision-making in critical societal infrastructure.
Praveen Kumar Donta (SM’22), Associate Professor (Docent) at the Department of Computer and Systems Sciences, Stockholm University, Sweden. Earlier, he was a Postdoc at Distributed Systems Group, TU Wien, Vienna, Austria. He received his Ph. D. at the Indian Institute of Technology (Indian School of Mines), Dhanbad, from the Department of Computer Science and Engineering in May 2021. He was a visiting PhD student at the University of Tartu, Estonia, during his PhD for six months. He received his Bachelor’s and Master’s in Technology from JNT University, Anantapur, India. He is an Editorial Board member for IEEE IoT Journal, Nature Scientific Reports, Computing Springer, and ETT Wiley. His current research includes the Learning techniques for Distributed Computing Continuum Systems.
Sindri Magnússon is an Associate Professor in the Department of Computer and Systems Science at Stockholm University, Sweden. He received a B.Sc. degree in Mathematics from the University of Iceland, Reykjav´ık Iceland, in 2011, a Masters degree in Applied Mathematics (Optimization and Systems Theory) from KTH Royal Institute of Technology, Stockholm, Sweden, in 2013, and the PhD in Electrical Engineering from the same institution, in 2017. He was a postdoctoral researcher 2018-2019 at Harvard University, Cambridge, MA, and a visiting PhD student at Harvard University in 2015 and 2016. His research interests include large-scale distributed/parallel optimization, machine learning, and control.